# An Efficient Algorithm for Generating Minimal Unique-Cause MC/DC Test cases for Singular Boolean Expressions

### v3

@@ Abstract.
Modified Condition/Decision Coverage (MC/DC) is a mandatory structural coverage criterion for ensuring the reliability and safety of critical systems. While its strictest form, Unique-Cause MC/DC, offers the highest assurance, research on its efficient test generation has been lacking. This gap is particularly significant, as an analysis of large-scale avionics systems shows that 99.7% of all conditional decisions are, in fact, Singular Boolean Expressions (SBEs) the ideal structure for applying Unique-Cause MC/DC. This paper proposes ‘Robin’s Rule’, a deterministic algorithm that directly constructs a minimal test set of $N+1$ cases to guarantee 100% Unique-Cause MC/DC for SBEs with $N$ conditions, without generating a full truth table. To validate our approach, we constructed a benchmark by reformulating the TCAS-II specifications into SBEs and verified the results using an industry-standard, certified commercial tool. The results confirm that our method consistently achieves 100% coverage with the theoretical minimum number of tests and is more efficient than the commercial tool. This work provides a practical and provably optimal solution for verifying safety-critical systems, ensuring both rigor and efficiency.
@@ 1. Introduction
    Experimental Proof of Correctness and Optimality: We empirically prove the correctness and optimality of the proposed algorithm. Through extensive experiments on a benchmark dataset based on the real-world safety-critical TCAS-II specification, we show that the generated $N+1$ test cases consistently achieve 100% Unique-Cause MC/DC coverage without exception when validated by a DO-178C certified commercial tool.
@@
The remainder of this paper is organized as follows. Section 2 discusses background and related work. Section 3 details the proposed algorithm. Section 4 presents the experimental evaluation, and Section 5 discusses the results. Finally, Section 6 concludes the paper.
@@ 2.1.2. Formal Definitions of Unique-Cause and Masking MC/DC
2.1.3. SBE (Singular Boolean Expressions) and Coupled Conditions
@@ 2.2. Related Work
The automated generation of test cases for MC/DC is a well established research area that has been actively studied over the past decades (Cao2024, ), (Shekhawat2021, ), (Barisal2021, ), (Godboley2017, ), (Kangoye2015, ), (Haque2014, ), (Ghani2009, ), (Awedikian2009, ). Early and theoretical approaches included using the Boolean Difference Function or the Independence Graph model to theoretically prove the feasibility of constructing a minimal $N+1$ test set. Modern research has built upon this foundation, proposing more sophisticated and automated generation techniques.
@@ 3.2. Phase 2: Temporary Table Generation
Based on the preprocessed SBE, an empty table is created to store the final test cases. For an SBE with $N$ operands, a two-dimensional table of size $(N+1)\times N$ is initialized to generate the theoretical minimum of $N+1$ test cases. Each row in this table corresponds to a single test case, and each column corresponds to a single condition. The specific procedure is described in Algorithm 2. It distinguishes between cases where the number of conditions is two versus more than two. If there are only two conditions, it proceeds directly to test case generation based on the operator (AND/OR).
@@ 3.4. Phase 4: Test Case Generation
21:  if IsExceptionCase($T_{expanded}$then
@@ 3.4.3. Exception Handling
    To handle such exceptions, Robin’s Rule applies a predefined rule as shown in the example in Figure 4 of the source thesis. This figure illustrates the test table generation process for a complex SBE like $({B_{i}}^{\wedge}\vee{B^{\vee}}_{i+1})\wedge({B^{\vee}}_{i+2}\wedge S_{i})$.
@@ 4. Experimental Evaluation
To evaluate the correctness, minimality, and efficiency of the proposed algorithm, ‘Robin’s Rule’, we conducted an experiment designed to answer three key Research Questions (RQs). This chapter details our experimental setup, research questions, and an analysis of the results.
@@ 4.1.1. Benchmark Dataset
This experiment is based on the Traffic Collision Avoidance System II (TCAS-II) specification, a well-known and complex real-world benchmark for safety-critical systems (Weyuker1994, ), As the proposed algorithm is optimized for SBEs, we constructed our benchmark dataset by modifying the 15 core logical expressions from the original TCAS-II benchmark into SBE form. This transformation is based on the fact that the vast majority (99.7%) of conditional statements in large-scale avionics software are SBEs (Chilenski2001, ). A comprehensive benchmark suite with a number of conditions ($N$) ranging from 5 to 23 was configured, as detailed in Table 2.
Table 2. TCAS-II SBE specification
No SBE Condition
1 !(a && b) && (c && !d && !e || !f && g && !h || !i && !j && !k) && (l && m && (n || o) && p || q && (r || s) && !t || u && (v || w)) 23
2 a && (!b || !c) && d || e 5
3 a && (!b || !c || d && e && !(!f && g && h && !i || !j && k && l) && !(!m && n && o && p || !q && !r && s)) || t 20
4 (!a && b || c && !d) && !(e && f) && !(g && h) && !(i && j) && ((k && l || m && n) && o && (!p || !q && !r || !s && (!t || !u))) 21
5 (!a && b || c && !d) && !(e && f) && !(g && h) && ((i && j || k && l) && m && (n && o || !p && q)) 17
6 !(a && b) && (!c && d && !e && !f && (g && h || !i && j)) 10
7 a && !b && !c && d && !e && f && (g || !h && (i || j)) && !(k && l || !m && n || o) 15
8 a && !b && !c && !((d && (e || !f && (g || h))) || i && (j || !k && (l || m)) && !n && !o) && !(p && q || !r && s && !t) 20
9 a && !b && !c && (d && (e || !f && (g || h)) && (!i && !j || k ) || !l) && (m && n || !o && p && !q) 17
10 a || b || c || !d && !e && f && g && !h && !i || j && (k || l) && !m 13
11 a && b && (c || d) && e || f && (g || h) && !i || j && (k || l) 12
12 a && (( b || c || d) && e || f && g || h && (i || j || k || l)) || (m || n) && (o || p || q) && r 18
13 (a && b || c && d) && e && (f || (g && (h && i || j && k))) 11
14 (a && b || c && d) && e && (f && g || !h && i) 9
15 !a && b && !c && !d && (e && f || !g && h) 8
@@ 4.1.2. Experimental Procedure
    Source Code Generation: For each of the 15 SBEs in the benchmark dataset, a C language source file was automatically generated, containing the respective SBE within an ‘if’ statement.
@@
    Test Case Generation: Using the implemented Robin’s Rule tool, each C source file was analyzed to generate a set of test cases satisfying Unique-Cause MC/DC, which was then saved in CSV file format. This process was highly efficient, completing in under 5 seconds on average for all SBEs.
@@ 4.1.3. Validation Tool and Comparison Target
To serve as an oracle for verifying the correctness and coverage of the generated test cases, we selected VectorCAST (ver. 2024). VectorCAST is an industry-standard commercial software that has received Tool Qualification for avionics standards such as DO-178C (Santhanam2007, ), (DO330, ), (Godboley2021a, ). Therefore, its analysis results can serve as a highly reliable baseline. The primary goal of this experiment is to objectively prove the ‘correctness’ of the proposed algorithm, rather than to compete with other tools on performance. The comparison of the number of test cases with VectorCAST is presented as a strong example to demonstrate that the ‘efficiency’ of our algorithm is at a practical level.
@@ 4.2. Research Questions (RQs)
    (Correctness): Do the $N+1$ test cases generated by our algorithm achieve 100% Unique-Cause MC/DC coverage when verified by a certified tool?
@@
    (Efficiency): How does the size of the test set generated by Robin’s Rule compare to that generated by an industry-standard commercial tool?
@@
This section describes how Robin’s Rule generates a minimal test set, using the benchmark with the most conditions, SBE No. 1 (N=23), as a target.
@@ 4.3.2. Test Case Generation Results
Table 3. The 24 Test Cases Generated by Robin’s Rule for SBE No. 1 (N=23).
Test Case a b c d e f g h i j k l m n o p q r s t u v w
@@ 4.4. Threats to Validity
This threat concerns potential flaws within our own experimental setup. There could be an error in our implementation of the Robin’s Rule algorithm. We mitigated this threat by cross-validating our results: for every SBE, the test set generated by our tool was independently verified by an external, certified tool (VectorCAST) to achieve 100% coverage, which provides strong evidence for the correctness of our implementation.
@@ 4.5. Overall Result and Analysys
To answer RQ1, we analyzed the number of test cases generated by Robin’s Rule. For every SBE specification with $N$ conditions, our algorithm generated exactly $N+1$ test cases. This result was consistent across the entire benchmark, from N=5 to N=23, empirically confirming that Robin’s Rule successfully generates a test set of the theoretical minimum size for Unique-Cause MC/DC.
@@
To answer RQ2, each of the $N+1$ test sets was validated using VectorCAST. The results show that 100% Unique-Cause MC/DC coverage was achieved in all 15 cases. This demonstrates the correctness and effectiveness of the proposed deterministic generation logic.
Answering RQ3: Comparison with Commercial Tool
To address RQ3, we compared the number of test cases generated by our algorithm to the results from VectorCAST’s own test case generator. Robin’s Rule consistently generated a smaller test set in every case. The reduction in test cases ranged from one to six, with an average reduction of 2.87 test cases per SBE. This clearly demonstrates the practical efficiency and optimality achieved by the proposed deterministic approach compared to existing industry-grade tools.
In summary, our experiment provides strong evidence that Robin’s Rule is a correct, effective, and efficient method for generating a minimal test set for Unique-Cause MC/DC for SBEs. The detailed data for these results can be found in Table 4.
Table 4. Test Generation Results: Robin’s Rule vs. VectorCAST.
SBE No. of VectorCAST Robin’s Rule
No. Cond. ($N$) Cases Coverage Cases Coverage
1 23 27 100% 24 100%
2 5 8 100% 6 100%
3 20 26 100% 21 100%
4 21 28 100% 22 100%
5 17 22 100% 18 100%
6 10 13 100% 11 100%
7 15 17 100% 16 100%
8 20 24 100% 21 100%
9 17 20 100% 18 100%
10 13 16 100% 14 100%
11 12 15 100% 13 100%
12 18 22 100% 19 100%
13 11 15 100% 12 100%
14 9 12 100% 10 100%
15 8 10 100% 9 100%
5. Discussion

### v4

@@ Abstract.
Modified Condition/Decision Coverage (MC/DC) is a mandatory structural coverage criterion for assuring the reliability of safety-critical software. Among its variants, Unique-Cause MC/DC provides the strongest assurance, yet efficient and scalable test generation for Unique-Cause MC/DC remains underexplored. This gap is particularly important because large-scale avionics studies report that 99.7% of conditional decisions are Singular Boolean Expressions (SBEs), for which Unique-Cause obligations can be precisely characterized. We propose Robin’s Rule, a deterministic, direct-construction algorithm that generates a provably minimal test suite of $N+1$ test cases to guarantee 100% Unique-Cause MC/DC for SBEs with $N$ conditions, without enumerating the $2^{N}$ truth table. The algorithm runs in $O(N^{2})$ time by explicitly constructing an $(N+1)\times N$ test table. To evaluate the approach, we build a benchmark of 25 SBEs consisting of 15 TCAS-II–derived conditions and 10 randomly generated SBEs with diverse operators and nesting. We validate achieved coverage using VectorCAST as an oracle and compare against state-of-the-art baseline paradigms using BDD and SAT. Across all benchmarks, Robin’s Rule consistently achieves 100% Unique-Cause MC/DC with the theoretical minimum number of tests, while providing stable and efficient generation time. This work offers a practical and provably optimal solution for Unique-Cause MC/DC test generation on SBEs, improving both rigor and scalability for safety-critical verification.
@@ 1. Introduction
    Experimental Demonstration of Correctness and Optimality: We empirically evaluate the correctness and optimality of the proposed algorithm. Through extensive experiments on a benchmark dataset based on the real-world safety-critical TCAS-II specification, we show that the generated $N+1$ test cases consistently achieve 100% Unique-Cause MC/DC coverage across all SBEs when validated using VectorCAST as an independent coverage oracle.
@@
The remainder of this paper is organized as follows. Section 2 discusses background and related work. Section 3 details the proposed algorithm. Section 4 presents the experimental evaluation and discusses the results. Finally, Section 5 concludes the paper and outlines future work.
@@ 2.1.2. Formal Definitions of Unique-Cause and Masking MC/DC
2.1.3. Singular Boolean Expressions (SBE) and Coupled Conditions
@@ 2.2. Related Work
The automated generation of test cases for MC/DC is a well-established research area that has been actively studied over the past decades (Cao2024, ), (Shekhawat2021, ), (Barisal2021, ), (Godboley2017, ), (Kangoye2015, ), (Haque2014, ), (Ghani2009, ), (Awedikian2009, ). Early and theoretical approaches included using the Boolean Difference Function or the Independence Graph model to theoretically prove the feasibility of constructing a minimal $N+1$ test set. Modern research has built upon this foundation, proposing more sophisticated and automated generation techniques.
@@ 3.2. Phase 2: Temporary Table Generation
Based on the preprocessed SBE, we initialize an empty table to store the final test cases. For an SBE with $N$ operands, we allocate a two-dimensional table of size $(N+1)\times N$, where each row corresponds to a single test case and each column corresponds to a condition. Algorithm 2 shows that the table is allocated with exactly $N+1$ rows (or $3$ rows when $N=2$, i.e., $N+1$), and subsequent phases only populate entries without changing the number of rows. Thus, Robin’s Rule outputs exactly $N+1$ test cases. It distinguishes between cases where the number of conditions is two versus more than two. If there are only two conditions, it proceeds directly to test case generation based on the operator (AND/OR).
@@
Lemma 1 (Test-suite cardinality). For any input SBE with $N\geq 2$ operands, Robin’s Rule generates exactly $N+1$ test cases.
[…]
@@ 3.4. Phase 4: Test Case Generation
21:  if IsExceptionCase($T_{expanded}$) then
@@ 3.4.3. Exception Handling
    To handle such exceptions, Robin’s Rule applies a predefined rule illustrated in Figure 4. This figure illustrates the test table generation process for a complex SBE like $({B_{i}}^{\wedge}\vee{B^{\vee}}_{i+1})\wedge({B^{\vee}}_{i+2}\wedge S_{i})$.
@@
3.5. Complexity Analysis
Let $N$ be the number of operands (conditions) in an input SBE. Robin’s Rule deterministically constructs an (N+1)(N+1)×NN test table, and its overall runtime is dominated by the cost of filling this table.
Phase 1 (Preprocessing) consists of applying De Morgan’s laws recursively to push NOT operators down to operands and then performing structural sorting. The De Morgan transformation visits each operator/operand in the expression tree at most a constant number of times; therefore, it is linear in the expression size. Structural sorting establishes a deterministic processing order among blocks; even if sorting is implemented using a comparison-based method, it is at most O($N$ log $N$) with respect to the number of operands/blocks and is asymptotically dominated by later table-construction costs.
Phase 2 (Temporary Table Generation) initializes a 2D table with $(N+1)$ rows and $N$ columns, which requires O($N^{2}$) space allocation.
Phase 3 (Structural Analysis of SBE) decomposes the normalized SBE into B-forms and S-forms and generates the Relation Table (RT). Decomposition traverses the expression structure once to identify forms, and RT generation scans the resulting operator hierarchy to produce a truth-value sequence. Both operations are linear in the number of operands/blocks, i.e., O($N$).
Phase 4 (Test Case Generation) performs iterative table composition according to Algorithm 3. At each composition step, the algorithm expands and fills a portion of the (N+1)(N+1)×NN table by (i) laying out the current partial table with the next block’s pattern and (ii) filling the empty regions using RT. Across the entire execution, each cell of the final (N+1)(N+1)×NN table is assigned a truth value a constant number of times (once when it is first filled, and possibly once more in exceptional-case handling). Therefore, the cumulative cost of FillEmptyRegions/FillSColumn/HandleException is proportional to the number of cells, i.e., O($N^{2}$).
In summary, the overall time complexity is O($N^{2}$), and the space complexity is O($N^{2}$), mainly due to the explicit construction of the (N+1)(N+1)×NN test table. This polynomial complexity stands in contrast to truth-table-based approaches requiring $2^{N}$ combinations.
@@ 4. Experimental Evaluation
To evaluate the correctness, minimality, and efficiency of the proposed algorithm, ‘Robin’s Rule’, we conducted an experiment designed to answer three key Research Questions (RQs). This section details our experimental setup, research questions, and an analysis of the results.
@@ 4.1.1. Benchmark Dataset
This experiment is based on the Traffic Collision Avoidance System II (TCAS-II) specification, a well-known and complex real-world benchmark for safety-critical systems (Weyuker1994, ). As the proposed algorithm is optimized for SBEs, we constructed our benchmark dataset by modifying the 15 core logical expressions from the original TCAS-II benchmark into SBE form, and adding 10 randomly generated SBE conditions (with random N and random usage of &&, ——, !, and parentheses) to broaden diversity. This transformation is based on the fact that the vast majority (99.7%) of conditional statements in large-scale avionics software are SBEs (Chilenski2001, ). A comprehensive benchmark suite with a number of conditions ($N$) ranging from 5 to 26 was configured, as detailed in Table 2. SBEs No. 1-15 are reformulated from TCAS-II, while No. 16-25 are synthetically generated SBEs to introduce additional structural variety. The full textual forms of all 25 SBE expressions are provided in Appendix A to improve readability while preserving reproducibility.
Table 2. SBE benchmark specifications (full expressions are provided in Appendix A).
No Origin Condition (N) Full SBE
1 TCAS-II 23 App. A, Listing 1
2 TCAS-II 5 App. A, Listing 2
3 TCAS-II 20 App. A, Listing 3
4 TCAS-II 21 App. A, Listing 4
5 TCAS-II 17 App. A, Listing 5
6 TCAS-II 10 App. A, Listing 6
7 TCAS-II 15 App. A, Listing 7
8 TCAS-II 20 App. A, Listing 8
9 TCAS-II 17 App. A, Listing 9
10 TCAS-II 13 App. A, Listing 10
11 TCAS-II 12 App. A, Listing 11
12 TCAS-II 18 App. A, Listing 12
13 TCAS-II 11 App. A, Listing 13
14 TCAS-II 9 App. A, Listing 14
15 TCAS-II 8 App. A, Listing 15
16 Random 19 App. A, Listing 16
17 Random 11 App. A, Listing 17
18 Random 24 App. A, Listing 18
19 Random 18 App. A, Listing 19
20 Random 26 App. A, Listing 20
21 Random 13 App. A, Listing 21
22 Random 17 App. A, Listing 22
23 Random 7 App. A, Listing 23
24 Random 11 App. A, Listing 24
25 Random 18 App. A, Listing 25
@@ 4.1.2. Experimental Procedure
    Source Code Generation: For each of the 25 SBEs in the benchmark dataset, a C language source file was automatically generated, containing the respective SBE within an ‘if’ statement.
@@
    Test Case Generation: Using the implemented Robin’s Rule tool, each C source file was analyzed to generate a set of test cases satisfying Unique-Cause MC/DC, which was then saved in CSV file format.
@@ 4.1.3. Validation Tool and Comparison Target
To validate the correctness and coverage of the generated test cases, we selected VectorCAST (ver. 2025) as an oracle. VectorCAST is an industry-standard commercial tool qualified for avionics standards such as DO-178C (Santhanam2007, ; DO330, ; Godboley2021a, ). Therefore, its analysis results can serve as a reliable oracle for validating the achieved Unique-Cause MC/DC coverage.
In addition to this coverage-oracle validation, we compare our approach against two representative baseline paradigms used in state-of-the-art MC/DC test generation for automated Boolean reasoning in MC/DC test generation: (i) a BDD(roBDD)-based generator, following the roBDD-based MC/DC generation line (Ahishakiye2021, ), and (ii) a SAT-solver-based generator, following SAT-based MC/DC test generation approaches (Kitamura2018, ). We use Z3 as the backend solver for the SAT baseline due to its maturity and widespread adoption in satisfiability-based program analysis.
These baselines are complementary: BDD-based methods construct a canonical decision-graph representation that supports systematic extraction of distinguishing input assignments, while SAT-based methods encode coverage obligations as satisfiability constraints and leverage solver technology to compute satisfying assignments. Together, they provide strong and informative references for assessing whether our direct-construction approach offers advantages in minimality, efficiency, and coverage guarantees.
We report (i) the number of generated test cases, (ii) generation time, and (iii) achieved Unique-Cause MC/DC coverage across all 25 SBEs (Tables 46).
@@ 4.2. Research Questions (RQs)
    (Correctness): Do the $N+1$ test cases generated by our algorithm achieve 100% Unique-Cause MC/DC coverage when verified using VectorCAST as an independent coverage oracle?
@@
    (Comparison with Baselines): How does Robin’s Rule compare to BDD and SAT-based baselines in terms of (a)test suite size, (b) generation time, and (c) achieved Unique-Cause MC/DC coverage?
@@
This section describes how Robin’s Rule generates a minimal test set, using a representative large-scale benchmark, SBE No. 1 (N=23), as a target.
@@ 4.3.2. Test Case Generation Results
Table 3. The 24 Test Cases Generated by Robin’s Rule for SBE No. 1 (NN=23).
Test Case n o m l p r s q !t v w u c !d !e !f g !h !i !j !k !a !b
@@ 4.4. Threats to Validity
This threat concerns potential flaws within our own experimental setup. There could be an error in our implementation of the Robin’s Rule algorithm. We mitigated this threat by cross-validating our results: for every SBE, the test set generated by our tool was independently verified using VectorCAST as an external commercial coverage-analysis tool to achieve 100% coverage, which provides strong evidence for the correctness of our implementation.
@@ 4.5. Overall Results
To answer RQ1, we analyzed the number of test cases generated by Robin’s Rule. For every SBE specification with $N$ conditions, our algorithm generated exactly $N+1$ test cases. This result was consistent across the entire benchmark, from NN=5 to NN=26, empirically confirming that Robin’s Rule successfully generates a test set of the theoretical minimum size for Unique-Cause MC/DC.
@@
To answer RQ2, each of the $N+1$ test sets was validated using VectorCAST. The results show that 100% Unique-Cause MC/DC coverage was achieved in all 25 cases. This demonstrates the correctness and effectiveness of the proposed deterministic generation logic.
Answering RQ3: Comparison with BDD, SAT solver
To address RQ3, we compare Robin’s Rule against BDD(roBDD) and SAT(Z3) baselines. Table 4-6 summarize the numerical results, and Figures 6-8 visualize the same outcomes.
  • Test suite size (Table 4 and Figure 6): Robin’s Rule consistently generates exactly $N+1$ test cases for every SBE, while BDD and SAT often require more than $N+1$.
  • Generation time (Table 5 and Figure 7): SAT and Robin’s Rule remains stable as NN grows, whereas BDD shows a clear increasing trend in generation time with larger NN.
  • Coverage (Table 6 and Figure 8): Robin’s Rule achieves 100% Unique-Cause MC/DC coverage for all SBEs, while the baselines do not always reach full coverage.
Table 4 (Test suite size). Table 4 reports, for each of the 25 SBEs, the number of conditions $N$ and the resulting test-suite size produced by each generator. Robin’s Rule generates exactly N+1N{+}1 test cases for all SBEs (ranging from 6 to 27 cases), matching the theoretical minimum. In contrast, the BDD baseline matches N+1N{+}1 only in 11 out of 25 SBEs and otherwise requires additional tests (up to +7+7 over N+1N{+}1, with an average overhead of +1.68+1.68). The SAT baseline exceeds N+1N{+}1 in all 25 cases, requiring on average +6.6+6.6 additional tests and reaching the maximum overhead of +16+16 (SBE No. 4).
Table 4. Test Case Generation Results (Number of Test Cases)
SBE No. of BDD SAT Robin’s Rule
No. Cond. ($N$) Cases Cases Cases
1 23 26 38 24
2 5 7 8 6
3 20 27 27 21
4 21 22 38 22
5 17 18 32 18
6 10 11 16 11
7 15 16 18 16
8 20 23 26 21
9 17 20 27 18
10 13 15 17 14
11 12 13 15 13
12 18 19 23 19
13 11 12 18 12
14 9 10 16 10
15 8 9 13 9
16 19 22 24 20
17 11 12 13 12
18 24 28 33 25
19 18 26 27 19
20 26 34 42 27
21 13 17 18 14
22 17 21 26 18
23 7 8 9 8
24 11 13 20 12
25 18 21 29 19
Figure 6 (Test suite size vs. $N$). Figure 6 visualizes the Table 4 results as a function of $N$ (from N=5N{=}5 to N=26N{=}26). The Robin’s Rule curve overlaps the N+1N{+}1 lower bound across the entire range, indicating consistent minimality regardless of expression size. The BDD curve stays relatively close to the lower bound for smaller $N$ but diverges for larger and structurally complex SBEs, whereas the SAT curve remains consistently above the bound, reflecting that the SAT-based construction in our baseline typically produces additional assignments beyond the minimum.
Figure 6. Number of generated test cases (BDD vs SAT vs Robin’s Rule)
Table 5 (Generation time). Table 5 reports the wall-clock time (seconds) required to generate test cases for each SBE. The SAT baseline is the fastest overall (median 0.0073 s; range 0.0023-0.0247 s). Robin’s Rule remains stable and efficient (median 0.0565 s; range 0.0280-0.1326 s). In contrast, the BDD baseline exhibits a heavy-tailed distribution (median 2.468 s; range 0.0003-2448.046 s), including extreme outliers such as SBE No. 20 (N=26N{=}26), which requires 2448.046 s.
Table 5. Test Case Generation Time (s)
SBE BDD SAT Robin’s Rule
No. Time (s) Time (s) Time (s)
1 215.9581 0.0247 0.092557
2 0.0003 0.0023 0.028010
3 22.0323 0.0091 0.062134
4 54.5546 0.01 0.076227
5 2.6119 0.0074 0.053130
6 0.01397 0.0039 0.034554
7 0.541 0.0059 0.054029
8 24.531 0.0092 0.071639
9 2.6163 0.0076 0.098991
10 0.1128 0.0056 0.049013
11 0.054 0.0047 0.041849
12 5.009 0.0077 0.056544
13 0.0265 0.0044 0.057589
14 0.0052 0.003 0.039289
15 0.0025 0.0028 0.034258
16 13.1345 0.0083 0.074857
17 0.0287 0.0043 0.048177
18 434.7779 0.0131 0.091852
19 6.6683 0.0076 0.089465
20 2448.046 0.0135 0.132568
21 0.1241 0.0053 0.051261
22 2.468 0.0073 0.050424
23 0.0011 0.0026 0.034107
24 0.0263 0.0041 0.062772
25 5.7416 0.0074 0.067168
Figure 7 (Time vs. $N$). Figure 7 plots the same timing results, using a logarithmic scale to accommodate the large variance of BDD. As $N$ increases, SAT and Robin’s Rule stay within a narrow band (sub-second and stable), while BDD shows both a clear upward trend and high variance. This highlights the practical sensitivity of BDD-based generation to the expression structure and the resulting diagram size, even within the SBE class.
Figure 7. Test case generation time
Table 6 (Achieved Unique-Cause MC/DC coverage). Table 6 reports the achieved Unique-Cause MC/DC coverage validated by VectorCAST. Robin’s Rule achieves 100% coverage for all 25 SBEs. The SAT baseline reaches 100% in 23 out of 25 SBEs but drops to 27% for SBE No. 1 and 79% for SBE No. 18. The BDD baseline reaches 100% in 13 out of 25 SBEs, with the minimum coverage being 33% (SBE No. 1). These failures are consistent with masking effects caused by short-circuit semantics and the fact that solver/graph-based baselines do not uniformly enforce Unique-Cause obligations under a fixed implementation and configuration. Overall, the results indicate that, under our baseline implementations and settings, the solver/graph-based approaches do not provide a uniform guarantee of Unique-Cause MC/DC across all SBEs, whereas Robin’s Rule does.
Table 6. Unique-Cause MC/DC Coverage Results (%)
SBE BDD SAT Robin’s Rule
No. Coverage (%) Coverage (%) Coverage (%)
1 33% 27% 100%
2 60% 100% 100%
3 75% 100% 100%
4 100% 100% 100%
5 100% 100% 100%
6 100% 100% 100%
7 100% 100% 100%
8 90% 100% 100%
9 94% 100% 100%
10 92% 100% 100%
11 100% 100% 100%
12 100% 100% 100%
13 100% 100% 100%
14 100% 100% 100%
15 100% 100% 100%
16 100% 100% 100%
17 90% 100% 100%
18 79% 79% 100%
19 100% 100% 100%
20 65% 100% 100%
21 92% 100% 100%
22 100% 100% 100%
23 100% 100% 100%
24 72% 100% 100%
25 72% 100% 100%
Figure 8 (Coverage visualization). Figure 8 summarizes Table 6 and highlights the key outcome: Robin’s Rule is the only approach that consistently attains full (100%) Unique-Cause MC/DC coverage across all 25 SBEs, while the baselines exhibit non-trivial coverage gaps on specific SBEs.
Figure 8. MC/DC test results
Overall, these results indicate that Robin’s Rule simultaneously ensures minimality $(N+1)$, stable efficiency, and guaranteed 100% Unique-Cause MC/DC coverage across the benchmark.
4.6. Discussion